OpenStudy (anonymous):

find the integral (e^x) / ((e^x)-1) dx

zepdrix (zepdrix):

\[\large \int\limits \frac{e^x\;dx}{e^x-1}\] Apply a `U-substitution`. Let your `u` be the denominator.

OpenStudy (anonymous):

if u=e^x -1 the du/dx= e^x

zepdrix (zepdrix):

Good, now move the dx to the other side c: giving you, du= e^x dx

zepdrix (zepdrix):

Understand how you'll plug those pieces in? :D

OpenStudy (anonymous):

kind of... would it be...the integral of e^x du/dx?

OpenStudy (anonymous):

no you get e^x du = e^x e^x dx

OpenStudy (anonymous):

is that right? or am i getting it confused?

zepdrix (zepdrix):

Ah sorry, I got distracted.

zepdrix (zepdrix):

\[\large \color{royalblue}{u=e^x-1} \qquad \qquad \color{salmon}{du=e^x \; dx}\] Just switch the blue with blue, and the pink with pink to get them plugged in correctly. \[\large \int\limits\limits \frac{\color{salmon}{e^x\;dx}}{\color{royalblue}{e^x-1}}\]

zepdrix (zepdrix):

\[\large \int\limits\limits \frac{\color{salmon}{e^x\;dx}}{\color{royalblue}{e^x-1}} \qquad \rightarrow \qquad \int\limits\limits \frac{\color{salmon}{du}}{\color{royalblue}{u}}\]

zepdrix (zepdrix):

does that make sense? :)

OpenStudy (anonymous):

up to that point yes....but then do im confused as to how to continue